# COAT: Compressing Optimizer states and Activation for Memory-Efficient FP8 Training

### v1

@@ 2 Related Work
Activations consume a significant amount of memory during training Cai et al. (2020), with memory usage scaling in proportion to sequence length and batch size. ActNN (Chen et al., 2021) is a memory-efficient training framework designed to reduce the memory footprint by quantizing activations to 2 bits. Few-bit Backward (Novikov et al., 2023) focus on quantizing activation functions while achieving higher performance. Jetfire proposes INT8 data flow to quantize the activation of both linear and non-linear layers to reduce memory footprint.
@@ 3 Preliminaries
To perform optimizer state quantization, we adopt per-group quantization for both first-order and second-order momentum, following previous works (Dettmers et al., 2021; Li et al., 2024). Every consecutive $G$ element forms a group ($G$ is defined as the group size), and each group is quantized independently with its own statistics. Optimizer states are stored in FP8 precision, while its scaling factor is stored in BF16. More details can be found in Appendix A.
@@ 4.2 Dynamic Range Expansion
where $k$ is used to control the strength of the expansion.
@@ 5.1 Decompose the activation memory footprint
In the forward pass of neural networks, activations must be preserved for the backward pass to calculate gradients. As illustrated in Table 2, non-linear layers such as LayerNorm/RMSNorm  (Ba, 2016; Zhang & Sennrich, 2019) and Activation Functions  (Hendrycks & Gimpel, 2016; Shazeer, 2020) typically account for approximately 50% of the memory footprint in the Llama model series Touvron et al. (2023). In contrast, linear layers contribute less than 25%. Therefore, it is essential to optimize both linear and non-linear layers to reduce activation memory footprint. [^2] RoPE and FlashAttention is more as a distinct module and should be handled separately.
@@ 6.1.1 LLM pretraining
To evaluate our method when pretraining LLMs, We train OLMo-1B (Groeneveld et al., 2024) and OLMo-7B on Dolma (Soldaini et al., 2024). Following the official report, we use a global batch size of 4M tokens (2048 macro batch size, with a sequence length of 2048 tokens). We use PyTorch FSDP in our experiments. For OLMo-1B, we pretrain for 5500 steps, which corresponds to approximately 22B tokens. For OLMo-7B, due to the resource limitation, we perform continue pretraining for 1000 steps (about 4B tokens) and resume from the official checkpoint at 5000 steps.
We report the training curve in Figure  5, and report the perplexity and accuracy result in Table 3. The training curve and downstream task performance were consistent with BF16 training and TransformerEngine baseline, validating the effectiveness of COAT. For OLMo-7B experiment, we report the training curve in Figure 6, where COAT also aligns well with the baseline and is nearly lossless.
Table 3: OLMo-1B pretraining performance on downstream tasks. TE refers to TransformerEngine.
@@
BF16 60.0% 45.6% 67.3% 33.7% 51.6%
TE 62.0% 45.4% 63.9% 33.8% 51.3%
COAT 61.0% 44.2% 67.6% 33.7% 51.5%
Figure 5: OLMo-1B pretraining loss curve.
Figure 6: OLMo-7B training curve.
@@ 6.1.2 LLM fine-tuning
We further evaluate our method on LLM fine-tuning task. We focus on math corpus and fine-tune a Llama-2-7B model on the MAmmoTH (Yue et al., 2023) dataset. We train for 3 epochs, and report the downstream tasks performance in  4. After fine-tuning, COAT still performs consistently with the baselines on the downstream task performance, proving the accurateness of our method.
@@ Appendix D Detailed Explanation for Table 2
Therefore, after FP8 quantization, the memory usage of linear layers comes from: QKV Projection - 0.5U, Attention Projection - 1U, Up&Gate Projection - 0.5U, Down Projection - 1.33U. They sum up to 3.33U. Our FP8 precision flow method is also compatible with other quantized attention method Shah et al. (2024); Zhang et al. (2024) by directly quantize the output of the RoPE function.

### v2

@@ 2 Related Work
Recent research has focused on reducing memory footprint Cai et al. (2020) during neural network training through activation quantization. ActNN (Chen et al., 2021) introduced a 2-bit activation compressed training framework using random quantization, achieving 12x reduction in activation memory footprint1. GACT (Liu et al., 2022a) extended this concept to support various machine learning tasks and architectures, providing up to 8.1x reduction in activation memory for CNNs, transformers, and GNNs. However, these methods mostly focus on convolutional networks and are not applicable to LLMs. AQ-SGD (Wang et al., 2022) proposed compressing activation changes rather than direct values for pipeline-parallel training to reduce the communication overhead in the pipeline-parallel setting. Few-Bit Backward(Novikov et al., 2023) quantizes the non-linear activation functions using optimal piecewise-constant approximations to reduce the memory footprint while maintaining convergence. Jetfire (Xi et al., 2024) proposes INT8 data flow to quantize the activation of both linear and non-linear layers to reduce memory footprint and is applicable to language model pretraining.
Bit-Width Under-Utilization
Balance Quantization Zhou et al. (2017) recursively partitions the parameters by percentiles to reduce quantization error. DDSQ Wang & Kang (2023) dynamically changes the quantization parameters according to different gradient distributions. N2UQ Liu et al. (2022b) improves the quantization precision via learning input thresholds and nonuniform mapping. These approaches are aware of the dynamic range problem in quantization and propose more adaptive, distribution-aware techniques that can preserve the nuanced information contained in full precision. They share a fundamental goal of expanding the effective dynamic range of low-precision neural network representations. These methods adopt a learnable non-uniform quantization lookup table, while our work relies on the natural FP8 data format.
@@ 3 Preliminaries
To perform optimizer state quantization, we adopt per-group quantization for both first-order and second-order momentum, following previous works (Dettmers et al., 2021; Li et al., 2024). Every consecutive $G$ element forms a group ($G$ is defined as the group size), and each group is quantized independently with its own statistics. Optimizer states are stored in FP8 precision, while its scaling factor is stored in BF16. In FSDP or ZeRO-3, the statistics related to optimizer states (e.g. scaling factor) are not synchronized across GPUs, since each GPU maintains its own shard of optimizer states. More details can be found in Appendix A.
@@ 4.2 Dynamic Range Expansion
where $k$ is used to control the strength of the expansion. In Appendix. F we prove that it is optimal.
@@ 5.1 Decompose the activation memory footprint
In the forward pass of neural networks, activations must be preserved for the backward pass to calculate gradients. As illustrated in Table 2, non-linear layers such as LayerNorm/RMSNorm  (Ba, 2016; Zhang & Sennrich, 2019) and Activation Functions  (Hendrycks & Gimpel, 2016; Shazeer, 2020) typically account for approximately 50% of the memory footprint in the Llama model series Touvron et al. (2023). In contrast, linear layers contribute less than 25%. Therefore, it is essential to optimize both linear and non-linear layers to reduce activation memory footprint. [^2] RoPEs and FlashAttentions are more as a distinct module and should be handled separately.
@@ 6.1.1 LLM pretraining
To evaluate our method when pretraining LLMs, We train OLMo-1B (Groeneveld et al., 2024) and OLMo-7B on Dolma (Soldaini et al., 2024). Following the official report, we use a global batch size of 4M tokens (2048 macro batch size, with a sequence length of 2048 tokens). We use PyTorch FSDP in our experiments.
For OLMo-1B, we conduct pretraining for 300B tokens, which corresponds to 75k training steps. We report the training curve in Figure  5, and report the perplexity and accuracy result in Table 3. The training curve and downstream task performance were consistent with BF16 training baseline, validating the effectiveness of COAT. For OLMo-7B experiment, we pretrain for 160B tokens, which corresponds to 40k training steps.we report the training curve in Figure 6 and downstream task performance in Table 4, where COAT also aligns well with the baseline and is nearly lossless.
Table 3: OLMo-1B pretraining performance on downstream tasks. We report the performance after training for 300B tokens.
@@
BF16 77.0 % 57.3 % 84.0 % 54.5 % 68.2 %
COAT 75.0 % 58.1 % 83.9 % 54.3 % 67.8 %
Figure 5: OLMo-1B training loss curve.
Table 4: OLMo-7B pretraining performance on downstream tasks. We report the performance after training for 160B tokens.
Train Loss WikiText C4 Pile Avg ppl
BF16 2.458 13.378 13.911 9.260 12.183
COAT 2.446 13.466 13.466 9.356 12.266
COPA ARC(Easy) SciQ HellaSwag Avg Acc
BF16 75.0% 61.2% 76.1% 60.0% 68.1 %
COAT 73.0% 63.0% 78.3% 59.7% 68.5 %
Figure 6: OLMo-7B training loss curve.
@@ 6.1.2 LLM fine-tuning
We further evaluate our method on LLM fine-tuning task. We focus on math corpus and fine-tune a Llama-2-7B model on the MAmmoTH (Yue et al., 2023) dataset. We train for 3 epochs, and report the downstream tasks performance in  5. After fine-tuning, COAT still performs consistently with the baselines on the downstream task performance, proving the accurateness of our method.
@@ Appendix A Details about optimizer states quantization
In the FSDP setting, dynamic range synchronization across GPUs is unnecessary due to the unique distributed training architecture. The approach leverages sharding strategies that eliminate the need for cross-GPU synchronization. Each GPU independently handles its own optimizer state shards. Since dynamic range is calculated based on local device statistics and optimizer states remain locally computed and updated, we do not need to synchronize the dynamic range across GPUs.
@@ Appendix D Detailed Explanation for Table 2
Therefore, after FP8 quantization, the memory usage of linear layers comes from: QKV Projection - 0.5U, Attention Projection - 1U, Up&Gate Projection - 0.5U, Down Projection - 1.33U. They sum up to 3.33U.
@@ Appendix E Detailed Explanation for Figure 1(c)
Appendix F Proof the expand function is optimal
We prove that the choice of $f(x)=x^{k}$ is optimal given certain basic assumptions for the expansion function $f(x)$.
  1. 1.
    $f(x)$ is an odd function, where $f(x)+f(-x)=0$, since the representation range of FP8 is symmetric. Therefore, in the following discussion, we only consider the case when $x\geq 0$.
  2. 2.
    $f(x)$ is continuously differentiable.
  3. 3.
    $f(x)$ is monotonically increasing. This is because we want to keep the relative order before and after quantization or expansion.
  4. 4.
    $f(0)=0$ and $f(1)=1$. $f(0)=0$ since $f(x)$ is odd. $f(1)=1$ since we can always consider $g(x)=\frac{f(x)}{f(1)}$ without changing its expansion ability.
  5. 5.
    For any $x>y>0,r>0$, $\frac{f(x)}{f(y)}=\frac{f(rx)}{f(ry)}$. This is used to make sure the function is scale-invariant. For example, if you multiply the unquantized value $X$ by $10$, the value after expansion and quantization is the same as if you do not multiply the value by $10$.
We prove that based on these assumptions, the only function that meets these assumptions is $f(x)=\text{sign}(x)|x|^{k}$.
Proof: We only consider the case when $x>0$. We denote $f(2)=m>1$. From assumption 5, we have for any $p>r>s>q$, if $pq=rs$, then $f(p)f(q)=f(r)f(s)$, since
$\frac{f(p)}{f(r)}=\frac{f\big{(}p\cdot(s/p)\big{)}}{f\big{(}r\cdot(s/p)\big{)}}=\frac{f(s)}{f(q)}.$
And for any $p=rs$,
$f(p)=f(p)f(1)=f(r)f(s).$
Therefore, for any $x=2^{t},t\in\mathbb{N}$,
$f(x)=f(2^{t})=f(2)^{t}=m^{t}.$
It is also true that for any $x=2^{-t},t\in\mathbb{N}$,
$f(x)=\frac{f(1)}{f(\frac{1}{x})}=\frac{f(1)}{f(2^{t})}=\frac{1}{m^{t}}=m^{-t},\text{so that }f(x)=m^{-t}.$
Therefore, for any $N\in\mathbb{N}$, $x=2^{t}$, $t=\sum\limits_{i=-N}^{N}c_{i}\cdot 2^{i}$, $c_{i}\in\{0,1\}$,
$f(x)=f(\prod_{i=-N}^{N}2^{c_{i}\cdot 2^{i}})=\prod_{i=-N}^{N}f(c_{i}\cdot 2^{i})=\prod_{i=-N}^{N}m^{c_{i}\cdot 2^{i}}=m^{\sum\limits_{i=-N}^{N}c_{i}\cdot 2^{i}}=m^{t}.$
For $\forall x>0$, it can be written in the form of $x=2^{t},t\in\mathbb{R}$. Consider the binary representation of $t$, denoted by $t=\sum\limits_{i=-\infty}^{\infty}c_{i}\cdot 2^{i}$, where $c_{i}\in\{0,1\}$. For any integer $N>0$, there exists two series $\{a_{i}\}_{-N}^{N}$ and $\{b_{i}\}_{-N}^{N}$ that satisfies $a_{-N}=0,b_{-N}=1,\text{and }a_{i}=b_{i}\text{ for }-N+1\leq i\leq N,$
$t_{a}\leq t<t_{b},$
where $t_{a}=\sum\limits_{i=-N}^{N}a_{i}\cdot 2^{i}$ and $t_{b}=\sum\limits_{i=-N}^{N}b_{i}\cdot 2^{i}$. Since $f(x)$ is continuous, the values of $f(x)$ at $x_{a}=2^{t_{a}}$ and $x_{b}=2^{t_{b}}$ can be used to approximate $f(x)=f(2^{t})$, and by taking the limit as $N\to\infty$, the equality holds:
$f(x)=\lim_{N\to\infty}f(x_{a})=\lim_{N\to\infty}f(x_{b}).$
Therefore for $\forall~{}t\in\mathbb{R}$, $f(2^{t})=m^{t}$ for some $m>0$. Therefore $\forall~{}x>0$, $f(x)=m^{\log_{2}(x)}$. Using the exponential and logarithmic identity $m^{\log_{2}(x)}=x^{\log_{2}(m)}$, this simplifies to:
$f(x)=x^{\log_{2}(m)},$
where $\log_{2}(m)$ is a constant determined by $f(2)=m$. If we denote $\log_{2}(m)=k$, then $f(x)=x^{k}$.
Appendix G Implementation Details of COAT
This section highlights the key differences and details of the implementation of the COAT LLaMA model. The primary focus is on the techniques used to handle quantization and optimization efficiently.
G.1 Model Architecture
The COAT LLaMA implementation introduces three new modules, which are not present in the Hugging Face implementation, specifically designed to optimize for FP8 precision flow:
  1. 1.
    BeforeAttention module (inherit from torch.autograd.Function) calculates the LayerNorm and QKV Projection before RoPE and correctly handles how quantized inputs interact with the residual connection. If Attention is quantized (Zhang et al., 2024b; a; Shah et al., 2024), then the output of this module should also be quantized to accommodate the input of Attention.
  2. 2.
    AfterAttention (inherit from torch.autograd.Function) calculate the attention projection layer and ensures quantized outputs are correctly scaled and summed up with residual connection.
  3. 3.
    MLPResidual (inherit from torch.autograd.Function) handles quantization within the feed-forward network (FFN) efficiently. It also incorporates residual connections in this module to handle it efficiently.
These modules ensure that the computation flow after quantization is correct. We fuse the group scaling method into the add operator in residual connection, therefore we need to manually control how the gradient flows in the network.
G.2 Triton-Based FP8 Kernels for Linear Layers
The linear layers utilize Triton to implement custom FP8 matrix multiplication kernels. This implementation introduces the following optimizations:
  1. 1.
    Block Size Tuning: The block sizes for dimensions $M$, $N$, and $K$ are autotuned, with options chosen from {128, 256}. This dynamic tuning optimizes performance for the specific hardware environment.
  2. 2.
    Group Scaling Technique: The output of linear layer should be BF16 or FP8 per-group quantized tensor. If it should be per-group quantized, to implement it efficiently in Triton, we reshape the output (BLOCK_M, BLOCK_N) to (BLOCK_M, BLOCK_N // G, G), where G is the group size. We then calculate the maximum value along the last dimension to get the scaling factor.
G.3 Triton-Based FP8 Kernels for Non-Linear Layers
  1. 1.
    Transposed FP8 Output: Besides generating the quantized FP8 output, these layers also produce a transposed FP8 output. This is critical for backward computations, where column-major inputs are required to calculate weight gradients.
  2. 2.
    Block Size Tuning: Block sizes for dimensions $M$ and $N$ are autotuned, with options chosen from {32, 64}.
G.4 Optimizer States
We implement a fused kernel for optimizer states update in CUDA. Since optimizer states use $1\times 128$ quantization group size, every CUDA block calculate the result of a quantization group and contains 128 threads, each thread corresponds to one element. We uses warp shuffle functions for reduction to get the maximum absolute value and minimum absolute value within a quantization group to perform dynamic range expansion. We
Direct computation of $X_{\text{min}}^{k}$ can result in underflow and numerical instability since its magnitude is usually about $10^{-12}$ and $k>3$. To mitigate this issue, we introduce an additional FP32 value $C_{X}$, defined as $C_{X}=\sqrt{X_{\text{min}}\cdot X_{\text{max}}}$, to ensure numerical stability. By normalizing $X$ with $C_{X}$, the values $\frac{X_{\text{min}}}{C_{X}}$ and $\frac{X_{\text{max}}}{C_{X}}$ are computed, which are reciprocals of each other and closer to 1 compared to the original $X_{\text{min}}^{k}$. This normalization enhances numerical robustness and reduces the risk of instability.
G.5 Others
  1. 1.
    Backward Gradient Handling: Backward gradients are computed in BF16 precision and are not quantized. This ensures that the gradient computations remain stable and accurate, avoiding numerical issues that can arise with quantization. Therefore for the backward kernels, the activation stored from forward is in FP8, while the input gradient and the output gradient are both in BF16 precision.
  2. 2.
    Gradient Accumulation Optimization: During the first gradient accumulation step, the scaling factors of weight matrices are stored. This avoids the need to recompute scaling factors in subsequent steps, reducing computational overhead. The FP8 weights are not stored to prevent additional memory usage, which is critical for large-scale models.
Appendix H Group Size Analysis
For activation, we calculate the quantization error of the output of last layer in forward pass, and the gradient before the first layer in backward pass, comparing with BF16. Based on Table below, we find that group size = 16 achieves a good balance for memory efficiency and quantization error.
Only Quantize Linear Also Quantize Activation
MSE 1 $\times$ 128 Group Per Tensor 4 8 16 32 64
Memory Overhead - - 50% 25% 12.5% 6.5% 3.2%
Forward 372 438 464 480 492 512 520
Backward 285 286 319 334 363 361 371
Table 10: Quantization overhead, forward, and backward MSE for different configurations and group sizes.
We also find that applying per-group quantization to linear layers is harmful for speed, while not improving the accuracy by too much. Even with system optimizations, this finer-grained method reduces the speed of linear layers by about 40% (when group size = 128, speed is reduced from 1300 TFlops to about 900 TFlops), greatly reducing the benefit of FP8. On the other hand, the accuracy gain is only marginal. Therefore we do apply per-tensor quantization for linear layers.
Appendix I Efficiency Breakdown for optimizer and activation
We break down the improvement of each component and present the result in this table. Quantizing the optimizer states to FP8 is helpful in reducing the memory footprint, and can potentially increase the batch size to fully utilize the GPU. Quantizing the activation to FP8 can accelerate the training, and this number can be further improved by increasing the maximum batch size.
Table 11: Llama-2-7B Training Memory and Throughput Analysis
Configuration Optimizer Activation Peak Max BS Throughput Speedup
Llama-2-7B on 2 GPU
BF16 26.2 GB 25.8 GB OOM 1 6130 tokens/s 1.00x
FP8 Optimizer 6.4 GB 25.8 GB 61.9 GB 2 7258 tokens/s 1.18x
FP8 Optimizer + FP8 Activation 6.4 GB 16.9 GB 35.6 GB 4 11351 tokens/s 1.85x
Llama-2-7B on 4 GPU
BF16 13.1 GB 25.8 GB 55.1 GB 2 7730 tokens/s 1.00x
FP8 Optimizer 3.2 GB 16.9 GB 44.7 GB 4 8384 tokens/s 1.08x
FP8 Optimizer + FP8 3.2 GB 16.9 GB 35.6 GB 4 11257 tokens/s 1.45x
Llama-2-7B on 8 GPU
BF16 6.5 GB 25.8 GB 41.2 GB 4 8238 tokens/s 1.00x
FP8 Optimizer 1.6 GB 16.9 GB 36.1 GB 4 8444 tokens/s 1.02x
FP8 Optimizer + FP8 Activation 1.6 GB 16.9 GB 27.0 GB 8 11241 tokens/s 1.36x
Appendix J Overhead introduced by Group Scaling
We show that the max reduction on each 1×G “can be seamlessly fused with the previous operation, adding minimal overhead”. Take RMSNorm as an example, the speed of the RMSNorm kernel only increases by less than 5% when quantization is fused with it. Therefore, this adds only minimal overhead to the kernel.
Table 12: Performance Comparison: With and Without Quantization
Matrix Size With Quantize Without Quantize
(4096, 1024) 10.1 ms 8.7 ms
(4096, 2048) 29.6 ms 28.7 ms
(4096, 4096) 55.6 ms 53.2 ms
(4096, 8192) 122.6 ms 117.5 ms
Appendix K Understand how each component influences the model performance
We separately quantize the optimizer and activations to better understand how each component influences the model performance. We find that quantizing activation incurs almost no degradation due to our well-designed quantization flow, and the main degradation comes from optimizer quantization. We choose to quantize both to maximize memory saving.
Train PPL COPA ARC (Easy) SciQ HellaSwag Average
BF16 2.995 60.0 45.6 67.3 33.7 51.6
FP8 O + A 3.008 61.0 44.2 67.6 33.7 51.5
FP8 O 2.998 (+0.003) 60.0 (+0.0) 44.5 66.0 34.1 51.1
FP8 A 2.999 (+0.004) 63.0 (+3.0) 44.0 68.4 34.1 52.3
Table 13: Performance comparison across different metrics and configurations.